# Error Analysis of Bayesian Inverse Problems with Generative Priors

### v1

@@ 1 Introduction
Hence, the design and choice of the prior $\mu$ is of great interest in the theory, application, and algorithmic development of Bayesian inverse problems (BIPs). Traditionally, the priors are designed by experts by balancing theoretical and algorithmic convenience with the specific needs of the problem at hand; see for example [20] or [32, Sec. 3.3] as well as Section 1.2. However, in the past decade, thanks to the rise of machine learning (ML) and generative modeling, a new class of methodologies have been proposed where the prior $\mu$ is learned or designed in an adaptive manner. This class largely falls under the category of data-driven methods for solving inverse problems in the parlance of [5, 46, 44].
@@ 2.2 Perturbation bounds with respect to Wasserstein distances
By convention we set ${\mathcal{W}}_{p}(\gamma,\gamma^{\prime})=+\infty$ if either measure does not belong to $\mathbb{P}^{p}(\mathcal{U})$. By the celebrated Kantorovich duality [62, Thm. 5.10 and Case. 5.16] for the particular case $p=1$ we have the equivalent expression
@@
We are now ready to present our first main theoretical results which allows us to control the ${\mathcal{W}}_{1}$ distance between posteriors in terms of the ${\mathcal{W}}_{2}$ distance between the priors. Our result extends those given in [55, Sec. 5] and [19, Sec. 4.3] to the case of likelihoods that are not globally Lipschitz.
@@ 3.1 Prior bounds
Note that [18, Thm. 1] covers the case $d<4$ but this mostly leads to unwieldy expressions for us that have little bearing on the main implications of our theory since $d>4$ is reasonable for most practical inverse problems. In fact, henceforth we assume that q>2d/(d2)q>2d/(d-2) so that the term $N^{-(q-2)/q}$ can be dropped and we have the simpler rate $N^{-2/d}$. $\Diamond$
@@
Suppose $d>4$, $\mu,\eta\in\mathbb{P}(\mathbb{R}^{d})$, and $\chi_{q}(\mu),\chi_{q}(\widehat{T}^{\dagger}\#\eta)<+\infty$ for some q>2d/(d2)q>2d/(d-2). Then there exists a constant $C>0$, independent of $N$ and $M$, such that for any $\epsilon>0$ it holds with probability $(1-C\chi_{q}(\mu)N^{-1/d}/\epsilon)(1-C(\chi_{q}(\widehat{T}^{\dagger}\#\eta)M^{-1/d}/\epsilon)$ that
@@
${\mathcal{W}}_{2}(\widehat{\mu}^{N,M},\mu)={\mathcal{W}}_{2}(\widehat{T}^{N,M}\#\eta^{M},\mu)$
$\leq{\mathcal{W}}_{2}(\widehat{T}^{N,M}\#\eta^{M},\mu^{N})+{\mathcal{W}}_{2}(\mu^{N},\mu)\quad\text{(By the Triangle Inequality)}$
$\leq{\mathcal{W}}_{2}(\widehat{T}^{\dagger}\#\eta^{M},\mu^{N})+{\mathcal{W}}_{2}(\mu^{N},\mu)\quad\text{(By optimality of $\widehat{T}^{N,M}$)}$
$\leq{\mathcal{W}}_{2}(\widehat{T}^{\dagger}\#\eta^{M},\widehat{T}^{\dagger}\#\eta)+{\mathcal{W}}_{2}(\widehat{T}^{\dagger}\#\eta,\mu^{N})+{\mathcal{W}}_{2}(\mu^{N},\mu)\quad\text{(By the Triangle Inequality)}$
$\leq{\mathcal{W}}_{2}(\widehat{T}^{\dagger}\#\eta^{M},\widehat{T}^{\dagger}\#\eta)+{\mathcal{W}}_{2}(\widehat{T}^{\dagger}\#\eta,\mu)+2{\mathcal{W}}_{2}(\mu^{N},\mu)\quad\text{(By the Triangle Inequality)}$
Notice that the end result is primarily different from (15) due to the term ${\mathcal{W}}_{2}(\widehat{T}^{\dagger}\#\eta^{M},\widehat{T}^{\dagger}\#\eta)$ which is precisely the empirical error of the generated prior with the best approximation of $T^{\dagger}$ in $\widehat{\mathcal{T}}$. The middle term can be controlled by (16) while we can apply Lemma 3.4 to the first and third terms to get,
$\mathbb{E}\>{\mathcal{W}}_{2}(\widehat{T}^{\dagger}\#\eta^{M},\widehat{T}^{\dagger}\#\eta)\leq C\chi_{q}(\widehat{T}^{\dagger}\#\eta)M^{-1/d},\qquad\mathbb{E}\>{\mathcal{W}}_{2}(\mu^{N},\mu)\leq C\chi_{q}(\mu)N^{-1/d}.$
@@
${\mathcal{W}}_{2}(\widehat{T}^{\dagger}\#\eta^{M},\widehat{T}^{\dagger}\#\eta)+W_{2}(\mu^{N},\mu)<\epsilon$
with probability $(1-2C\chi_{q}(\widehat{T}^{\dagger}\#\eta)M^{-1/d}/\epsilon)(1-2C\chi_{q}(\mu)N^{-1/d}/\epsilon)$. Absorbing the independent constants into $C$ completes the proof.
@@ 3.2.1 The bounded support case
Then there exists an independent constant $C>0$ so that for any $\epsilon>0$, with probability $(1-C\chi_{q}(\mu)N^{-1/d}/\epsilon)(1-C\chi_{q}(\widehat{T}^{\dagger}\#\eta)M^{-1/d}/\epsilon)$, it holds that
@@ 3.2.2 The unbounded support case
Further write $\widehat{\nu}^{N,M}_{r}$ for the posterior measure arising from $\widehat{\mu}^{N,M}_{r}$. Write $T^{\dagger}_{r}$ for the map that satisfies $\mu_{r}=T^{\dagger}_{r}\#\eta$ and define $\widehat{T}^{\dagger}_{r}:=\operatorname*{arg\,min}_{T\in\widehat{T}}\|T-T^{\dagger}_{r}\|_{L^{2}(\eta)}$.
@@
    $\chi_{q}(\mu),\chi_{q}(\widehat{T}^{\dagger}_{r}\#\eta)<+\infty$ for some q>2d/(d2)q>2d/(d-2).
Then there exists a constant $C>0$ so that for any $\epsilon>0$, with probability $(1-C\chi_{q}(\mu_{r})N^{-1/d}/\epsilon)(1-C\chi_{q}(\widehat{T}^{\dagger}_{r}\#\eta)M^{-1/d}/\epsilon)$, it holds that

### v2

@@ 1 Introduction
Hence, the design and choice of the prior $\mu$ is of great interest in the theory, application, and algorithmic development of Bayesian inverse problems (BIPs). Traditionally, the priors are designed by experts by balancing theoretical and algorithmic convenience with the specific needs of the problem at hand; see for example [21] or [33, Sec. 3.3] as well as Section 1.2. However, in the past decade, thanks to the rise of machine learning (ML) and generative modeling, a new class of methodologies have been proposed where the prior $\mu$ is learned or designed in an adaptive manner. This class largely falls under the category of data-driven methods for solving inverse problems in the parlance of [5, 47, 45].
@@ 2.2 Perturbation bounds with respect to Wasserstein distances
By convention we set ${\mathcal{W}}_{p}(\gamma,\gamma^{\prime})=+\infty$ if either measure does not belong to $\mathbb{P}^{p}(\mathcal{U})$. By the celebrated Kantorovich duality [63, Thm. 5.10 and Case. 5.16] for the particular case $p=1$ we have the equivalent expression
@@
We are now ready to present our first main theoretical results which allows us to control the ${\mathcal{W}}_{1}$ distance between posteriors in terms of the ${\mathcal{W}}_{2}$ distance between the priors. Our result extends those given in [56, Sec. 5] and [20, Sec. 4.3] to the case of likelihoods that are not globally Lipschitz.
@@ 3.1 Prior bounds
Note that [19, Thm. 1] covers the case $d<4$ but this mostly leads to unwieldy expressions for us that have little bearing on the main implications of our theory since $d>4$ is reasonable for most practical inverse problems. In fact, henceforth we assume that q>2d/(d2)q>2d/(d-2) so that the term $N^{-(q-2)/q}$ can be dropped and we have the simpler rate $N^{-2/d}$. $\Diamond$
@@
Lemma 3.9.
Suppose $T\in L^{2}(\eta)$ is globally Lipschitz with constant $\ell_{T}$, then, for any pair of measures $\eta,\eta^{\prime}\in\mathbb{P}(\mathcal{U})$ it holds that ${\mathcal{W}}_{2}(T\#\eta,T\#\eta^{\prime})\leq\ell_{T}{\mathcal{W}}_{2}(\eta,\eta^{\prime})$.
Proof 3.10.
Let $\pi^{\star}$ denote the optimal ${\mathcal{W}}_{2}$ coupling between $\eta$ and $\eta^{\prime}$. Then using the Lipschitz assumption on $T$ we can write,
${\mathcal{W}}_{2}^{2}(T\#\eta,T\#\eta^{\prime})$ $=\inf_{\pi\in\Pi(\eta,\eta^{\prime})}\int_{\mathcal{U}\times\mathcal{U}}\|T(x)-T(y)\|_{\mathcal{U}}^{2}\;\pi({\rm d}x,{\rm d}y)$
$\leq\int_{\mathcal{U}\times\mathcal{U}}\|T(x)-T(y)\|_{\mathcal{U}}^{2}\;\pi^{\star}({\rm d}x,{\rm d}y)=\ell_{T}^{2}\int_{\mathcal{U}\times\mathcal{U}}\|x-y\|_{\mathcal{U}}^{2}\;\pi^{\star}({\rm d}x,{\rm d}y)$
@@
Suppose $d>4$, $\mu,\eta\in\mathbb{P}(\mathbb{R}^{d})$, and $\chi_{q}(\mu),\chi_{q}(\eta)<+\infty$ for some q>2d/(d2)q>2d/(d-2). Further suppose that the elements of $\widehat{\mathcal{T}}$ are uniformly globally Lipschitz with constant $\ell_{\widehat{\mathcal{T}}}>0$. Then there exists a constant $C>0$, independent of $N$ and $M$, such that for any $\epsilon>0$ it holds with probability $(1-\ell_{\widehat{\mathcal{T}}}C\chi_{q}(\eta)M^{-1/d}/\epsilon)(1-C\chi_{q}(\mu)N^{-1/d}/\epsilon)$ that
@@
${\mathcal{W}}_{2}(\widehat{\mu}^{N,M},\mu)={\mathcal{W}}_{2}(\widehat{T}^{N,M}\#\eta,\mu)$
$\leq{\mathcal{W}}_{2}(\widehat{T}^{N,M}\#\eta,\widehat{T}^{N,M}\#\eta^{M})+{\mathcal{W}}_{2}(\widehat{T}^{N,M}\#\eta^{M},\mu)\quad\text{(By the Triangle Inequality)}$
$\leq\ell_{\widehat{\mathcal{T}}}{\mathcal{W}}_{2}(\eta,\eta^{M})+{\mathcal{W}}_{2}(\widehat{T}^{N,M}\#\eta^{M},\mu^{N})+{\mathcal{W}}_{2}(\mu^{N},\mu)$
(By the Triangle Inequality, Lemma 3.9)
$\leq\ell_{\widehat{\mathcal{T}}}{\mathcal{W}}_{2}(\eta,\eta^{M})+{\mathcal{W}}_{2}(\widehat{T}^{\dagger}\#\eta^{M},\mu^{N})+{\mathcal{W}}_{2}(\mu^{N},\mu)\quad\text{(By optimality of $\widehat{T}^{N,M}$)}$
$\leq\ell_{\widehat{\mathcal{T}}}{\mathcal{W}}_{2}(\eta,\eta^{M})+{\mathcal{W}}_{2}(\widehat{T}^{\dagger}\#\eta^{M},\widehat{T}^{\dagger}\#\eta)+{\mathcal{W}}_{2}(\widehat{T}^{\dagger}\#\eta,\mu^{N})+{\mathcal{W}}_{2}(\mu^{N},\mu)$
(By the Triangle Inequality)
$\leq 2\ell_{\widehat{\mathcal{T}}}{\mathcal{W}}_{2}(\eta,\eta^{M})+{\mathcal{W}}_{2}(\widehat{T}^{\dagger}\#\eta,\mu)+2{\mathcal{W}}_{2}(\mu^{N},\mu)$
(By the Triangle Inequality and applying Lemma 3.9 again)
The second term can be controlled by (16) while we can apply Lemma 3.4 to the first and third terms and absorbing the lipschitz constant to get,
$\mathbb{E}\>{\mathcal{W}}_{2}(\eta,\eta^{M})$ $\leq C\chi_{q}(\eta)M^{-1/d}\qquad\mathbb{E}\>{\mathcal{W}}_{2}(\mu^{N},\mu)\leq C\chi_{q}(\mu)N^{-1/d},$
@@
$2\ell_{\widehat{\mathcal{T}}}{\mathcal{W}}_{2}(\eta,\eta^{M})+2{\mathcal{W}}_{2}(\mu^{N},\mu)<\epsilon$
with probability $(1-4\ell_{\widehat{\mathcal{T}}}C\chi_{q}(\eta)M^{-1/d}/\epsilon)(1-2C\chi_{q}(\mu)N^{-1/d}/\epsilon)$. Absorbing the independent constants into $C$ completes the proof.
Remark 3.13.
We note that the uniform Lipschitz assumption on the approximation class $\widehat{\mathcal{T}}$ may appear restrictive but it is feasible in practical applications by, for example, weight clipping strategies or regularization for neural nets [64, 11]. $\Diamond$
@@ 3.2.1 The bounded support case
Then there exists an independent constant $C>0$ so that for any $\epsilon>0$, with probability $(1-\ell_{\widehat{\mathcal{T}}}C\chi_{q}(\eta)M^{-1/d}/\epsilon)(1-C\chi_{q}(\mu)N^{-1/d}/\epsilon)$, it holds that
@@ 3.2.2 The unbounded support case
where the approximation class $\widehat{\mathcal{T}}$ is globally $\ell_{\widehat{\mathcal{T}}}-$Lipschitz. Further write $\widehat{\nu}^{N,M}_{r}$ for the posterior measure arising from $\widehat{\mu}^{N,M}_{r}$. Write $T^{\dagger}_{r}$ for the map that satisfies $\mu_{r}=T^{\dagger}_{r}\#\eta$ and define $\widehat{T}^{\dagger}_{r}:=\operatorname*{arg\,min}_{T\in\widehat{\mathcal{T}}}\|T-T^{\dagger}_{r}\|_{L^{2}(\eta)}$, $\widehat{T}_{r}^{\star}:=\operatorname*{arg\,max}_{T\in\widehat{\mathcal{T}}}{\mathcal{W}}_{2}(T\#\eta,T\#\eta^{M})$.
@@
    $\chi_{q}(\mu),\chi_{q}(\eta)<+\infty$ for some q>2d/(d2)q>2d/(d-2).
Then there exists a constant $C>0$ so that for any $\epsilon>0$, with probability $(1-\ell_{\widehat{\mathcal{T}}}C\chi_{q}(\eta)M^{-1/d}/\epsilon)(1-C\chi_{q}(\mu)N^{-1/d}/\epsilon)$, it holds that